OpenStudy (anonymous):

What is the solution to the equation? -4x +5(x-2)=-16+10

OpenStudy (anonymous):

use this http://www.mathpapa.com/algebra-calculator.html

OpenStudy (hyroko):

first, multiply the 5 out over the terms inside the parenthesis then combine like terms. then get X on one side by itself.

OpenStudy (anonymous):

I'll draw what they said out ^

OpenStudy (anonymous):

ok

OpenStudy (hyroko):

5(x-2) = 5x-10 -4x+5x = x x-10 = -16 +10 add 10 to both sides x= -16+20 x=4

OpenStudy (anonymous):